# SCOPE: Stochastic Cartographic Occupancy Prediction Engine for Uncertainty-Aware Dynamic Navigation

### v1

@@ V-C Computational Resource Utilization
While it is encouraging that our proposed SCOPE family of OGM predictors exhibits good prediction performance, we also want to evaluate the inference speed and model size of all tested OGM predictors. This is because mobile robots are resource-limited, and smaller model sizes and faster inference speeds mean robots have a faster reaction time to face and handle dangerous situations in complex dynamic scenarios.
V-C1 Baseline Comparison
@@
Table II summarizes the inference speed and model size of six predictors tested on an NVIDIA Jetson TX2 embedded computer equipped with a 256-core NVIDIA Pascal @ 1300MHz GPU and loaded with 8GB of memory. We can see that: 1) DeepTracking [6] has the smallest model size, our software-optimized SO-SCOPE model is the next smallest, our SCOPE and SCOPE++ models are about 4.9 times larger than SO-SCOPE, ConvLSTM [5] is about 1.4 times larger than these, and PhyDNet [7] is about 3 times larger than ConvLSTM. 2) Our proposed SCOPE family achieves faster inference than all other state-of-the-art baselines, especially our SO-SCOPE which can run at up to 35 FPS. These show that our proposed SCOPE family (i.e., SCOPE++, SCOPE, and SO-SCOPE) can have faster inference speed and reasonable model size, which is more hardware-friendly than other state-of-the-art OGM predictors.
TABLE II: Inference Speed and Model Size of different OGM predictors
Models ConvLSTM [5] PhyDNet [7] DeepTracking [6] SCOPE-NEMC SCOPE SCOPE++ SO-SCOPE
Inference Speed (FPS) 2.95 4.66 5.32 24.83 23.29 10.68 34.75
Model Size (MB) 12.44 37.17 0.95 8.84 8.84 8.85 1.80
@@
Figure 10: Inference speed and memory usage of SCOPE and SO-SCOPE predictors with different number of samples.
V-C2 SCOPE Family Detailed Comparison
@@
As can be seen from Fig. 10, the inference speed of the SCOPE predictor drops significantly as the number of samples increases, with the lowest inference speed being about 2 FPS. In contrast, the inference speed of the SO-SCOPE hardly slows down as the number of samples increases until 128 samples are generated, and even with 1,024 samples it can still achieve around 6 FPS (i.e., the speed bottleneck is generating large number of samples rather than network inference). This demonstrates the significant advantage of using knowledge distillation techniques to compress our SCOPE++ network, since smaller networks lead to faster inference speeds, and using prediction uncertainty statistics lookup table to provide uncertainty estimates for our SO-SCOPE predictor, which speeds up the time-consuming sample generation process. Note that we could parallelize OGM sampling using our SO-SCOPE approach since it is simply requires querying the lookup table, while SCOPE requires running the VAE model.
Figure 10, also shows that the memory usage of the SCOPE predictor increases significantly with the number of samples and can only generate up to 32 samples on an NVIDIA Jetson TX2 embedded computer due to the 8 GB memory limit. In contrast, the memory usage of the SO-SCOPE predictor barely increases with the number of samples until 1,024 samples are generated (i.e., the memory bottleneck is storing large number of samples). This demonstrates the significant advantage in memory reduction of using the prediction uncertainty statistics lookup table to provide uncertainty estimates for our SO-SCOPE predictor.
V-D Uncertainty Characterization
To characterize the uncertainty information of our SCOPE, SCOPE++, and SO-SCOPE predictors, we use the Shannon entropy [76] to measure the uncertainty. For an OGM ${\mathbf{o}}$, the Shannon entropy is given by
$H(\mathbf{o})$ $=\sum_{c\in\mathbf{o}}H(c)$ (7a)
$=-\sum_{c\in\mathbf{o}}p\log p+(1-p)\log(1-p),$ (7b)
where the first equality comes from the assumption that cells $c$ within an OGM are independent and the second equality comes from the definition of Shannon entropy for a Bernoulli distribution with parameter $p$ (since each cell can either be occupied, with probability $p$, or free, with probability $1-p$).
In our case, we have a predicted future map, $\hat{\mathbf{o}}$ and data about the distribution of occupancy values $\tilde{c}$, which comes from the predicted occupancy value $\hat{c}$ along with the mixture model (6) and learned parameters $\mathbf{\xi}$. In this case, the entropy becomes
@@ V-D Uncertainty Characterization
V-D1 Experiment Setup
Our characterization of the uncertainty has three primary goals. First, to show that the generated OGMs are consistent with reality. Second, to show that the SCOPE family predictors converge to a consistent distribution. Third, to show that the uncertainty in the overall OGM increases with the number of objects in the scene.
For the experiments, we use the OGM-Turtlebot data dataset and fix the prediction horizon to $\tau=5$. We use the procedure outlined in Fig. 5 to count the number of objects in each OGM. Then, we randomly select 20 input sequences for each number of objects (from 1 to 12) and generate a total of 1,024 OGM samples for each input sequence using each SCOPE method.
V-D2 Qualitative results
@@
Figure 11: A diverse set of prediction samples of our SCOPE++, SCOPE, and SO-SCOPE predictors tested on the OGM-Turtlebot2 dataset at the 5th prediction timestep. The black area is the free space and the white area is the occupied space. The red bounding boxes highlight variations across outputs for the same input.
V-D3 Quantitative results
One of the biggest differences of our proposed methods compared to state-of-the-art baselines is that SCOPE family can provide uncertainty estimates. The goal of these tests is to analyze the output distribution of our SCOPE, SCOPE++, and SO-SCOPE predictors.
@@ V-D3 Quantitative results
First, we show how the entropy of the final probabilistic OGM changes as the number of OGM samples increases, with the hypothesis that it will level off at some value well below that of the entropy of a uniform distribution. Figure 12a shows that the average entropy of SCOPE(SAMPLE) and SCOPE++(SAMPLE) predictors (i.e., computed via (8c)) starts to converge and plateau at a certain value after generating 128 samples, which is consistent with our hypothesis. Note that the entropy values for SCOPE, SCOPE++, and SO-SCOPE are flat lines because we use (8b) for calculation and no sampling is required. Furthermore, the differences among the three flat lines are small (i.e., no more than 0.004), which indicates that our uncertainty quantification for SCOPE++ is correct and can provide reasonable uncertainty estimates for our SO-SCOPE predictor. The differences of them are mainly caused by the differences in their predicted OGMs. Figure 13 shows the differences in cell counts for each bin in their predicted OGMs. It is consistent with the entropy differences of our SCOPE, SCOPE++, and SO-SCOPE predictors, where the difference between SCOPE and SCOPE++ is smaller than the difference between SO-SCOPE and SCOPE++ (i.e., SO-SCOPE has a smaller percentage of cell counts or lower entropy near bin 1 to bin 11).
Figure 13: The cell counts for each bin in SCOPE, SCOPE++, and SO-SCOPE predictors.
However, we can see that there is a gap between the entropy calculated by (8b) using the mean map $\bar{\mathbf{o}}$ and the entropy calculated by (8c) (i.e., SCOPE/SCOPE++ is higher than SCOPE(SAMPLE)/SCOPE++(SAMPLE)). We believe this gap stems from two factors. First, our assumption that cells $c$ within an OGM are independent will overestimate the uncertainty values because actually knowing the occupancy of one cell should reduce the uncertainty of neighboring cells (and our learned latent representation in the VAE is encoding these correlations). Second, the mean map $\bar{\mathbf{o}}$ is different individual samples $\tilde{\mathbf{o}}$, so the statistics differ slightly.
@@ VI Uncertainty-Aware Navigation
We next test the applicability of our proposed OGM predictors to practical mobile robot navigation in crowded dynamic scenes. To leverage the uncertainty information in the environmental future states and provide robust and reliable navigation behavior, we propose a general navigation control policy framework (shown in Fig. 14) based on costmaps, which are used in the move_base ROS navigation framework and can be integrated with most currently exiting control polices. Specifically, we use prediction and uncertainty costmaps to tell the robot the potentially dangerous areas ahead of it and how certain they are, which enables mobile robots to make safer nominal path plans, take proactive actions to avoid potential collisions and improve navigation capabilities. Note that each costmap grid cell of our proposed prediction and uncertainty costmaps has an initial constant cost, and we map each occupied grid cell of the prediction costmap and uncertainty costmap to a Gaussian obstacle value rather than a “lethal” obstacle value. This is because the predicted obstacles and uncertainty regions are not real obstacle spaces.
Figure 14: System architectures of the SCOPE-based and SO-SCOPE-based predictive uncertainty-aware navigation planners. The blue font emphasizes the difference between the SCOPE-based navigation framework and the SO-SCOPE-based navigation framework. The basic process of our proposed navigation framework is: first, the lidar data is also fed into our SCOPE or SO-SCOPE predictor to generate predicted OGM samples or lookup statistics from the prediction uncertainty statistics lookup table. Then, we can easily generate the prediction mean map and uncertainty map from these samples or the statistics. Finally, we create two new costmap layers, the prediction costmap layer and the uncertainty costmap layer, combine them into the master costmap, and obtain our SCOPE-based or SO-SCOPE-based predictive uncertainty aware planners.
@@ VI-A Baselines and Evaluation Metrics
We first use the model-based DWA [8] as the local planner to instantiate a SCOPE-based predictive uncertainty-aware planner (i.e., DWA-SCOPE-PU) to demonstrate how prediction and its uncertainty information improve robot navigation performance. Then, we use the learning-based DRL-VO [11] as the local planner to instantiate a SO-SCOPE-based predictive uncertainty-aware planner (i.e., DRL-VO-SO-SCOPE-PU) to demonstrate the software-optimized SO-SCOPE predictor is hardware friendly and our proposed SO-SCOPE-based navigation framework can be applied to any currently exiting control policies.
We test these two predictive uncertainty-aware control policies, along with four state-of-the-art control policies: model-based DWA planner [8], supervised-learning-based CNN [9] (ours), DRL-based A1-RC [10], and DRL-based DRL-VO [11] (ours), and two ablation prediction-aware control policies without uncertainty costmaps: DWA-DeepTracking-P and DWA-SCOPE-P. Note that we use these four state-of-the-art baseline policies directly from their papers [8, 9, 10, 11] without any retraining or parameter tuning.
To evaluate the performance of navigation control policies, we use the following four metrics from [9, 11]:
  • Success rate: the fraction of collision-free trials.
  • Average time: the average travel time of trials.
  • Average length: the average trajectory length of trials.
  • Average speed: the average speed during trials.
@@ VI-B Experiment Setup
Following the navigation evaluation settings in work [11], all control policies are tested by a Turtlebot2 robot with a maximum speed of 0.5 m/s, equipped with a Hokuyo UTM-30LX lidar and a ZED stereo camera, in the simulated Lobby Gazebo world with 15-45 pedestrians (sampling interval 10), as shown in Fig. 15a. The measurement range of Hokuyo lidar is set to $[0.1,30]$ m, its FOV is $270^{\circ}$, and its angular resolution is $0.25^{\circ}$. The depth range of the ZED camera is set to $[0.3,20]$ m, and its FOV is $90^{\circ}$.
In addition, we also deploy our proposed DWA-SCOPE-PU and DRL-VO-SO-SCOPE-PU control policies on a real Turtlebot2 robot, which has the same configuration as the simulated robot but uses an NVIDIA Jetson AVG Xavier embedded computer as its its main computing device. We conduct testing in an indoor hallway environment at Temple University during peak traffic periods between classes, the floor plan of which is shown in Fig. 15b. Note that considering the computational resources of the Turtlebot2 robot, we use the SCOPE predictor to only generate 8 predicted OGM samples at the 6th prediction time step (i.e., 0.6 s) for the DWA-SCOPE-PU control policy and it cannot combine with any learning-based control polices. In contrast, our software-optimized SO-SCOPE predictions do not suffer from such problems and can be combined with model-based or learning control strategies (e.g., with DRL-VO in our instantiation).
@@
Lobby world, 15 pedestrians CNN [9] - - - -
@@
DRL-VO-SO-SCOPE-PU 0.97 11.69 5.43 0.46
Lobby world, 25 pedestrians CNN [9] 0.80 19.31 6.16 0.32
@@
DRL-VO-SO-SCOPE-PU 0.93 11.95 5.52 0.46
Lobby world, 35 pedestrians CNN [9] 0.81 14.30 5.40 0.38
@@
DRL-VO-SO-SCOPE-PU 0.92 12.26 5.57 0.45
Lobby world, 45 pedestrians CNN [9] 0.79 16.65 5.62 0.34
@@
DRL-VO-SO-SCOPE-PU 0.83 12.24 5.56 0.45

### v2

@@ V-D Computational Resource Utilization
While it is encouraging that our proposed SCOPE family of OGM predictors exhibits good prediction performance, we also want to evaluate the inference speed, model size, and memory usage of all tested OGM predictors. This is because mobile robots are resource-limited, and smaller model sizes, less memory usage, and faster inference speeds mean robots have a faster reaction time to face and handle dangerous situations in complex dynamic scenarios.
V-D1 Baseline Comparison
@@
Table II provides a comprehensive benchmark of resource usage, summarizing the inference speed, model size, and memory usage of nine predictors tested on a Jetson TX2 embedded computer equipped with a 256-core NVIDIA Pascal @ 1300MHz GPU and loaded with 8GB of memory. We can see that: 1. our SCOPE series achieves much faster inference than all other state-of-the-art baselines, especially our SO-SCOPE which can run at up to 35 FPS (about 89.1 times faster than the slowest SAAConvLSTM [8]and TAAConvLSTM [8]), 2. while our SCOPE and SCOPE++ models have the third-smallest model size, our proposed software-optimized SO-SCOPE model has the second-smallest model size and is approximately 894.7 times smaller than the largest LOPR [9] model, and 3. our SCOPE series has the second smallest memory usage, which is about 7.6 times smaller than LOPR which has the largest memory usage of 5GB. These show that our SCOPE family can have real-time inference speed, reasonable model size, and small memory usage, which is more hardware-friendly than other state-of-the-art OGM predictors and can be deployed on resource-constrained robots.
@@
Figure 7: Inference speed and memory usage of SCOPE and SO-SCOPE predictors with a different number of samples.
V-D2 SCOPE Family Detailed Comparison
@@
As can be seen from Fig. 7, the inference speed of the SCOPE predictor drops significantly as the number of samples increases, with the lowest inference speed being about 2 FPS. In contrast, the inference speed of the SO-SCOPE hardly slows down as the number of samples increases until 128 samples are generated, and even with 1,024 samples it can still achieve around 6 FPS (i.e., the speed bottleneck is generating a large number of samples rather than network inference). This demonstrates the significant advantage of using knowledge distillation techniques to compress our SCOPE++ network, since smaller networks lead to faster inference speeds, and using prediction uncertainty statistics lookup table to provide uncertainty estimates for our SO-SCOPE predictor, which speeds up the time-consuming sample generation process. Note that we could parallelize OGM sampling using our SO-SCOPE approach since it simply requires querying the lookup table, while SCOPE requires running the VAE model.
Figure 7 also shows that the memory usage of the SCOPE predictor increases significantly with the number of samples and can only generate up to 32 samples on a Jetson TX2 device due to its 8 GB memory limit. In contrast, the memory usage of the SO-SCOPE predictor barely increases with the number of samples until 1,024 samples are generated (i.e., memory bottleneck is storing a large number of samples). It demonstrates the significant advantage of using the prediction uncertainty statistics lookup table to provide uncertainty estimates for our SO-SCOPE predictor in memory reduction.
V-E Uncertainty Characterization
To characterize the uncertainty information of our SCOPE family of predictors, we use the Shannon entropy [75] to measure the uncertainty. For an OGM ${\mathbf{o}}$, the Shannon entropy is given by $H(\mathbf{o})=-\sum_{c\in\mathbf{o}}p\log p+(1-p)\log(1-p)$, where the first equality comes from the assumption that cells $c$ within an OGM are independent and the second equality comes from the definition of Shannon entropy for a Bernoulli distribution with parameter $p$ (since each cell can either be occupied, with probability $p$, or free, with probability $1-p$). In our case, we have a predicted map, $\hat{\mathbf{o}}$ and data about the distribution of occupancy values $\tilde{c}$, which comes from the predicted occupancy value $\hat{c}$ along with the mixture model (6) and learned parameters $\mathbf{\xi}$. Then, the entropy becomes
@@ V-E Uncertainty Characterization
V-E1 Experiment Setup
Our characterization of the uncertainty has three primary goals: 1. show that the generated OGMs are consistent with reality, 2. show that the SCOPE series converge to a consistent distribution, and 3. show that the uncertainty in the overall OGM increases with the number of objects in the scene. For the experiments, we use the OGM-Turtlebot data dataset and fix the prediction horizon to $\tau=5$. We use the evaluation pipeline used to calculate OSPA Error on OGMs in [4] to count the number of objects in each OGM. Then, we randomly select 20 input sequences for each number of objects (from 1 to 12) and generate a total of 1,024 OGM samples for each input sequence using each SCOPE method.
V-E2 Qualitative results
@@
Figure 8: A diverse set of prediction samples of our SCOPE++, SCOPE, and SO-SCOPE predictors tested on the OGM-Turtlebot2 dataset at the 5th prediction timestep. The red bounding boxes highlight the regions of variation in the predicted samples for the same input in each stochastic algorithm (each row), which demonstrates the diversity of our proposed stochastic predictors.
V-E3 Quantitative results
One of the biggest differences between our proposed methods compared to state-of-the-art baselines is that the SCOPE family can provide uncertainty estimates. The goal of these tests is to analyze the output distribution of our SCOPE, SCOPE++, and SO-SCOPE predictors.
@@ V-E3 Quantitative results
First, we show how the entropy of the final probabilistic OGM changes as the number of OGM samples increases, with the hypothesis that it will level off at some value well below that of the entropy of a uniform distribution. Figure 9a shows that the average entropy of SCOPE(SAMPLE) and SCOPE++(SAMPLE) predictors (i.e., computed via (7c)) starts to converge and plateau at a certain value after generating 128 samples, which is consistent with our hypothesis. Note that the entropy values for SCOPE, SCOPE++, and SO-SCOPE are flat lines because we use (7b) for calculation and no sampling is required. Furthermore, the differences among the three flat lines are small (i.e., no more than 0.004), which indicates that our uncertainty quantification for SCOPE++ is correct and can provide reasonable uncertainty estimates for our SO-SCOPE predictor. The differences between them are mainly caused by the differences in their predicted OGMs. Figure 10 shows the differences in cell counts for each bin in their predicted OGMs. It is consistent with the entropy differences of our SCOPE, SCOPE++, and SO-SCOPE predictors, where the difference between SCOPE and SCOPE++ is smaller than the difference between SO-SCOPE and SCOPE++ (i.e., SO-SCOPE has a smaller percentage of cell counts or lower entropy near bin 1 to bin 11).
Figure 10: Cell counts for each bin in the SCOPE family of predictors.
However, we can see that there is a gap between the entropy calculated by (7b) using the mean map $\bar{\mathbf{o}}$ and the entropy calculated by (7c) (i.e., SCOPE/SCOPE++ is higher than SCOPE(SAMPLE)/SCOPE++(SAMPLE)). We believe this gap stems from two factors. First, our assumption that cells $c$ within an OGM are independent will overestimate the uncertainty values because actually knowing the occupancy of one cell should reduce the uncertainty of neighboring cells (and our learned latent representation in the VAE is encoding these correlations). Second, the mean map $\bar{\mathbf{o}}$ is different from individual samples $\tilde{\mathbf{o}}$, so the statistics differ slightly.
@@
Figure 11: System architectures of the SCOPE-based and SO-SCOPE-based predictive uncertainty-aware navigation planners. The blue font emphasizes the difference between the SCOPE-based navigation framework and the SO-SCOPE-based navigation framework. The basic process of our proposed navigation framework is as follows: first, the lidar data is also fed into our SCOPE or SO-SCOPE predictor to generate predicted OGM samples or lookup statistics from the prediction uncertainty statistics lookup table. Then, we can easily generate the prediction mean map and uncertainty map from these samples or the statistics. Finally, we create the prediction costmap layer and the uncertainty costmap layer, combine them into the master costmap, and obtain our SCOPE-based or SO-SCOPE-based predictive uncertainty-aware planners.
@@ VI Uncertainty-Aware Navigation
We next test the applicability of our SCOPE series to practical mobile robot navigation in crowded dynamic scenes. To leverage the uncertainty information in the environmental future states and provide robust and reliable navigation behavior, we propose a general predictive uncertainty-aware navigation framework (shown in Fig. 11) based on costmaps. Although this costmap-based framework loses the stochastic consistency of the entire navigation system, the biggest advantage is that it is used in the move_base ROS navigation framework and can be integrated with most currently existing control policies. Specifically, we use prediction and uncertainty costmaps to tell the robot the potentially dangerous areas ahead of it and how certain they are, which enables mobile robots to make safer nominal path plans, take proactive actions to avoid potential collisions, and improve navigation capabilities. To generate prediction and uncertainty costmaps from the output of our SCOPE family, we first binarize the predicted OGM and its uncertainty map using an occupancy threshold. Second, we initialize a constant cost value for each binarized prediction and uncertainty map grid cell and generate the initial prediction and uncertainty costmaps. Finally, we map each occupied grid cell of the prediction costmap and uncertainty costmap to a Gaussian obstacle value rather than a “lethal” obstacle value and obtain the final costmaps. This is because the predicted obstacles and uncertainty regions are not real obstacle spaces.
@@ VI-A Baselines and Evaluation Metrics
Since our proposed predictive uncertainty-aware navigation framework can be combined with different types of existing control policies, we define the following naming convention for predictive uncertainty-aware control strategy instances: [policy]/[predictor]/[P/PU], where [policy] is the control policy used, [predictor] is the OGM predictor used, P means using only the prediction map, and PU means using both the prediction map and its uncertainty map.
Following this naming convention, we first use the model-based DWA [10] as the local planner to instantiate a SCOPE-based predictive uncertainty-aware planner (i.e., DWA/SCOPE/PU) to demonstrate how prediction and its uncertainty information improve robot navigation performance. Then, we use the learning-based DRL-VO [13] as the local planner to instantiate a SO-SCOPE-based predictive uncertainty-aware planner (i.e., DRL-VO/SO-SCOPE/PU) to demonstrate the SO-SCOPE predictor is hardware friendly and our SO-SCOPE-based navigation framework can be applied to any currently existing control policies.
We test these two predictive uncertainty-aware control policies, along with four state-of-the-art control policies: model-based DWA planner [10], supervised-learning-based CNN [11], DRL-based A1-RC [12], and DRL-based DRL-VO [13], and two ablation prediction-aware control policies without uncertainty maps: DWA/DeepTracking/P and DWA/SCOPE/P. Note that we use these four state-of-the-art baseline policies directly from their papers [10, 11, 12, 13] without any retraining or parameter tuning. To evaluate the performance of navigation policies, we use the following four metrics from [11, 13]: success rate, average time, average length, and average speed.
@@ VI-B Experiment Setup
Following the navigation evaluation settings in [11, 13], we conduct a total of 100 trials, and all control policies are tested by a Turtlebot2 robot with a maximum speed of 0.5 m/s, equipped with a Hokuyo UTM-30LX lidar and a ZED stereo camera, in the simulated Lobby Gazebo world with 15-45 pedestrians (sampling interval 10), as shown in Fig. 12a. The measurement range of Hokuyo lidar is set to $[0.1,30]$ m, its FOV is $270^{\circ}$, and its angular resolution is $0.25^{\circ}$. The depth range of ZED camera is set to $[0.3,20]$ m, and its FOV is $90^{\circ}$.
In addition, we deploy our proposed DWA/SCOPE/PU and DRL-VO/SO-SCOPE/PU control policies on a real Turtlebot2 robot, which has the same configuration as the simulated robot but uses an NVIDIA Jetson AVG Xavier embedded computer as its main computing device. We conduct testing in an indoor hallway environment at Temple University during peak traffic periods between classes, the floor plan of which is shown in Fig. 12b. Note that considering the computational resources of Turtlebot2, we use the SCOPE predictor to only generate 8 predicted OGM samples at the 6th prediction time step (i.e., 0.6 s) for the DWA/SCOPE/PU control policy and it cannot combine with any learning-based control policies. In contrast, our SO-SCOPE predictions do not suffer from such problems and can be combined with model-based or learning control strategies (e.g., with DRL-VO in our instantiation).
@@
Lobby world, 15 pedestrians DWA [10] 0.94 12.21 5.09 0.42
CNN [11] - - - -
@@
DWA/DeepTracking/P 0.95 12.10 5.03 0.42
DWA/SCOPE/P 0.93 12.62 5.06 0.40
DWA/SCOPE/PU 0.98 12.79 5.06 0.40
DWA/SO-SCOPE/P 0.94 12.12 5.10 0.42
DWA/SO-SCOPE/PU 0.96 13.25 5.12 0.39
DRL-VO/SO-SCOPE/PU 0.97 11.69 5.43 0.46
Lobby world, 25 pedestrians DWA [10] 0.82 13.49 5.12 0.38
CNN [11] 0.80 19.31 6.16 0.32
@@
DWA/DeepTracking/P 0.87 12.67 5.06 0.40
DWA/SCOPE/P 0.90 13.53 5.08 0.37
DWA/SCOPE/PU 0.91 13.75 5.07 0.37
DWA/SO-SCOPE/P 0.90 13.23 5.13 0.39
DWA/SO-SCOPE/PU 0.92 14.26 5.19 0.36
DRL-VO/SO-SCOPE/PU 0.93 11.95 5.52 0.46
Lobby world, 35 pedestrians DWA [10] 0.82 14.18 5.15 0.36
CNN [11] 0.81 14.30 5.40 0.38
@@
DWA/DeepTracking/P 0.84 13.93 5.10 0.37
DWA/SCOPE/P 0.86 13.79 5.12 0.34
DWA/SCOPE/PU 0.89 14.90 5.12 0.34
DWA/SO-SCOPE/P 0.86 14.03 5.16 0.37
DWA/SO-SCOPE/PU 0.88 15.53 5.20 0.33
DRL-VO/SO-SCOPE/PU 0.92 12.26 5.57 0.45
Lobby world, 45 pedestrians DWA [10] 0.77 15.39 5.16 0.34
CNN [11] 0.79 16.65 5.62 0.34
@@
DWA/DeepTracking/P 0.78 15.23 5.14 0.34
DWA/SCOPE/P 0.79 14.84 5.14 0.35
DWA/SCOPE/PU 0.82 15.96 5.17 0.32
DWA/SO-SCOPE/P 0.78 14.89 5.21 0.35
DWA/SO-SCOPE/PU 0.80 16.29 5.27 0.32
DRL-VO/SO-SCOPE/PU 0.83 12.24 5.56 0.45